OpenStudy (linn99123):

Show all steps used to solve the problem. One point for steps and one point for solution. Be sure that you check your solution for correctness. 5n + 6 = -4 (Points : 2)

OpenStudy (linn99123):

@bibby

OpenStudy (texaschic101):

5n + 6 = -4 --- subtract 6 from both sides 5n + 6 - 6 = -4 - 6 -- combine like terms 5n = -10 -- divide both sides by 5 n = -2 check.. 5n + 6 = -4 5(-2) + 6 = -4 -10 + 6 = -4 -4 = -4 (correct) n = -2 any questions ?

OpenStudy (linn99123):

no i just could not explain it was hard thank you